MnO2 on ignition converts into Mn3O4. A sample of pyrolusite having 75% MnO2,20%inert impurities and rest water is ignited in air to constant weight. What is the percentage of Mn in the ignited sample 

Grade:

1 Answers

abhishek upadhyay
12 years ago

sivaji

apply POAC Mn atom

let there be 100g mixture

then wtg. of Mn= 75g now, inert impurities =20g  and xg Mn3O4 is formed

                                   MnO2 = Mn3O4

               for Mn atom  75/87  =3x/229                                                     ( M.W. of MnO2 = 87, M.W. of Mn3O4 =229)

then x=65.8

wt of mn in product = (75/87) * 55=47.41g

%tage of Mn in product = 47.41/(65.8+20)  * 100=55.26%